Discourses in Mathematics and Its Applications, No. 3 1994

ISBN 0-9630728-2-X

A Tribute to Ilya Bakelman

Edited by I. R. Bakelman, S. A. Fulling, and S. D. Taliaferro

Preface

The mathematical community was saddened by the death of Ilya J. Bakelman on August 30, 1992. He will be greatly missed by his family and by his colleagues and friends around the world. In October 1993 the Mathematics Department of Texas A&M University hosted a conference as a tribute to him. Speakers at that conference and other authors have kindly provided personal reminiscences and scholarly contributions to this memorial volume.